# Course Allocation with Credits via Stable Matching

### v1

@@ 4.1 Arbitrary course preferences
Let us first consider the problem of finding a pair-size-stable matching. In order to do so, we use the following procedure. First, we group courses by number of credits, from largest to smallest. Let $\{O(c_{j}):c_{j}\in C\}=\{N_{1},\dots,N_{R}\}$ for some $R\geq 1$, where $N_{k}>N_{k+1}$ ($11\leq k\leq R-$ 1). Let $C_{k}=\{c_{j}\in C:O(c_{j})=N_{k}\}$.
@@
0:  Course Allocation instance $I$
0:  return a pair-size-stable matching $M$ in $I$
@@
2:  for $k\in 1\dots R$ do
3:     while some student $s_{i}\in S$ satisfies $O(C_{M}(s_{i}))+N_{k}\leq T(s_{i})$ and $s_{i}$ has not applied to all feasible courses in $C_{k}$ on her list do
4:        $c_{j}:=$ $s_{i}$’s most preferred feasible course in $C_{k}$ that $s_{i}$ has not yet applied to;
@@
It is clear that Algorithm 1 finds a matching $M$. Furthermore, if there are downward-feasible constraints, $M$ is feasible, because every $s_{i}$ applies only to courses $c_{j}$ such that adding $(s_{i},c_{j})$ to $M$ does not violate downward-feasible constraints. With appropriate data structures, the total complexity of Algorithm 1 is $O(nm^{2})$, where $n=|S|$ and $m=|C|$. This is because for every student we consider each course in her preference list at most $m$ times, if there are downward-feasible constraints, as courses that were previously unfeasible might become feasible for some $s_{i}$ if she is rejected from some $c_{j}$. Meanwhile, when a student applies to a course, the course can perform all necessary steps in constant time, giving a total time complexity of $O(nm^{2})$, where $n=|S|$ and $m=|C|$. If there are no downward-feasible constraints then the time complexity is $O(nm)$. We now show that $M$ is pair-size-stable.
@@
We will argue by contradiction. Assume the matching $M$ found by Algorithm 1 is not pair-size-stable, and that there is a size-blocking pair $(s_{i},c_{j})$, such that $c_{j}\in C_{l}$ for some $l$. Thus, either $s_{i}$ has enough remaining credits to match with $c_{j}$ (and adding $(s_{i},c_{j})$ to $M$ results in a feasible matching) or $s_{i}$ prefers $c_{j}$ over every course $c_{k}$ from a subset of courses $D_{i}$ such that $s_{i}$ is assigned to every $c_{k}\in D_{i}$, $O(c_{j})\geq\sum_{c_{k}\in D_{i}}O(c_{k})$, and, if there are downward-feasible constraints, $(M\cup\{(s_{i},c_{j})\})\backslash D_{i}$ is feasible.
Assume that the first case holds. The course $c_{j}$ is in subset $C_{l}$, meaning that it was considered in the $l$-th round. As $s_{i}$ has enough credits for $c_{j}$ after the execution of the algorithm, it had enough credits during round $l$, because matches between students and courses cannot be broken, so after every round the number of remaining credits for $s_{i}$ stays the same or decreases. Thus, $s_{i}$ applied to $c_{j}$ during round $l$, so if they are not matched at the end of round $l$ then $c_{j}$ rejected $s_{i}$ for a better student. Hence, $(s_{i},c_{j})$ was not a size-blocking pair at the end of round $l$. Because all matches formed in round $l$ are part of the final pair-size-stable matching, $c_{j}$ prefers all of its students over $s_{i}$, so $(s_{i},c_{j})$ is not a size-blocking pair, a contradiction. Therefore, this case cannot occur.
Now, assume the second case holds, so $s_{i}$ prefers $c_{j}$ over every course $c_{k}$ from a subset of courses $D_{i}$ such that $s_{i}$ is assigned to every $c_{k}\in D_{i}$ and $O(c_{j})\geq\sum_{c_{k}\in D_{i}}O(c_{k})$. As $O(c_{j})\geq\sum_{c_{k}\in D_{i}}O(c_{k})$, $O(c_{j})\geq O(c_{k})$ $\forall c_{k}\in D_{i}$, so no $c_{k}\in D_{i}$ has a number of credits strictly greater than $O(c_{j})$. Additionally, no $c_{k}\in D_{i}$ has a number of credits equal to $O(c_{j})$, for if there were such a $c_{k}$ and $(s_{i},c_{j})$ was a size-blocking pair, then that would imply that in the round where $c_{j}$ and $c_{k}$ were considered, Deferred Acceptance did not find a stable matching, which is a contradiction. Therefore, $c_{j}$ has a number of credits strictly greater than every $c_{k}\in D_{i}$. However, this means that, during the algorithm’s execution, $c_{j}$ was considered in an earlier round than every $c_{k}\in D_{i}$, round $l$. During that round, $s_{i}$ had enough remaining credits for $c_{j}$, because, after removing every course from $D_{i}$, it has enough credits for $c_{j}$ after the execution of the algorithm. Using a similar argument to the first case, $(s_{i},c_{j})$ was not a size-blocking pair at the end of round $l$, meaning that it is not a size-blocking pair upon termination of the algorithm, a contradiction. Therefore, this case also cannot occur. By assuming either case holds, we arrive at a contradiction, which is caused by the assumption that there exists a size-blocking pair. Therefore, no such pair exists, and the matching is pair-size-stable. ∎
@@
Figure 4: Instance of Course Allocation. This instance is derived from a Hospitals/Residents with Sizes instance due to McDermid and Manlove mcdermid_keeping_2010 .
Determining whether a pair-stable matching exists is NP-complete, which can be seen by considering a connection between course allocation and the problem Hospitals/Residents with Sizes, a modification of Hospitals/Residents in which some residents need more than one space mcdermid_keeping_2010 . With respect to the stability definition given by McDermid and Manlove mcdermid_keeping_2010 , the problem of deciding whether a stable matching exists, given an instance of Hospitals/Residents with Sizes, is NP-complete. Moreover this result holds even if each resident has size 1 or 2, and each hospital has upper quota at most 2. By replacing residents by courses and hospitals by students, we can then obtain a simple reduction to course allocation where each student $s_{i}$ satisfies $T(s_{i})\leq 2$ and each course $c_{j}$ satisfies $O(c_{j})\in\{1,2\}$. Moreover, McDermid and Manlove’s stability definition corresponds to pair stability. The following result is therefore immediate.
Theorem 3 (mcdermid_keeping_2010 ).
ca-p-find is NP-hard, even if there are no downward-feasible constraints.
Similarly, determining the existence of a first-coalition-stable matching is NP-hard, and we show that by transforming from Hospitals/Residents with Sizes, though in this case a more complex reduction is required.
@@
ca-fc-find is NP-hard, even if there are no downward-feasible constraints.
@@
We prove this result by transforming from the problem of determining the existence of a stable matching in the Hospitals/Residents with Sizes (HRS) setting, which was shown to be NP-complete by McDermid and Manlove mcdermid_keeping_2010 . In the HRS instance $I$ that we are transforming from, each resident has size $1$ or $2$, each hospital has upper quota $2$, and the length of every preference list is at most $3$.
@@
The transformation proceeds as follows. Every resident $r_{j}$ in $I$ becomes a course $c_{j}$ in the Course Allocation instance $J$, with upper quota $1$, and number of credits $1$ or $2$, depending on the size of $r_{j}$. Every hospital $h_{i}$ becomes a student $s_{i}$, with $T(s_{i})=2$. We use this bijection to transform every preference list. For example, if resident $r_{j}$ has preference list $h_{i}\succ_{r_{j}}h_{k}\succ_{r_{j}}h_{l}$, now course $c_{j}$ has preference list $s_{i}\succ_{c_{j}}s_{k}\succ_{c_{j}}s_{l}$. Additionally, for every student $s_{i}$ we create a dummy course $c_{i}^{\prime}$, of size $1$, and append this course to the end of $s_{i}$’s preference list. The preference list of $c_{i}^{\prime}$ contains $s_{i}$ only.
@@
Assume that the HRS instance $I$ has a stable matching $M$. Then we build a first-coalition-stable matching $M^{\prime}$ in $J$ as follows. We take $M$ and swap residents and hospitals for courses and students, respectively. Then, for every student $s_{i}$ such that $C_{M^{\prime}}(s_{i})<2$, we add the pair $(s_{i},c_{i}^{\prime})$ to $M$. This matching is first-coalition-stable. Assume it is not, which implies the existence of a first-blocking coalition involving one student $s_{i}$. Let us go through each possibility:
@@
  • Student $s_{i}$ has only $1$ credit (which must be from course $c_{i}^{\prime}$, according to the construction of $M^{\prime}$), and forms a first-blocking coalition with a course $c_{j}$, with size $1$ or $2$. In this case, hospital $h_{i}$ would have no residents in $M$, and resident $r_{j}$ would prefer $h_{i}$ over her other partner, so $(r_{j},h_{i})$ would block $M$ in $I$, a contradiction.
@@
  • Student $s_{i}$ has $2$ credits, one of which comes from $c_{i}^{\prime}$ and the other from some $c_{k}$, and she forms a first-blocking coalition with course $c_{j}$. If $c_{j}$ has size $1$, then $s_{i}$ prefers $c_{j}$ over $c_{i}^{\prime}$. However, this implies that in $M$ hospital $h_{i}$ had one resident only, with capacity for resident $r_{j}$, which preferred $h_{i}$ over her assigned hospital, forming a blocking pair, a contradiction. If $c_{j}$ has size $2$, then $s_{i}$ prefers $c_{j}$ over $c_{k}$, which implies that in $I$ hospital $h_{i}$ prefers resident $r_{j}$ over its assigned resident $r_{k}$, and $r_{j}$ prefers $h_{i}$ over her assigned hospital, forming a blocking pair, a contradiction.
@@
  • Student $s_{i}$ has $2$ credits, which come from two courses of size $1$, namely $c_{p}$ and $c_{q}$, or one course of size $2$, namely $c_{p}$, and she forms a first-blocking coalition with either one course $c_{j}$, or with two courses, $c_{j}$ and $c_{k}$, such that $c_{j}\succ_{s_{i}}c_{k}$. Then, in $I$, hospital $h_{i}$ prefers resident $r_{j}$ over either resident $r_{p}$ or residents $r_{p}$ and $r_{q}$, and $r_{j}$ prefers $h_{i}$ over her assigned hospital. Hence, $(r_{j},h_{i})$ blocks $M$ in $I$, a contradiction.
@@
Similarly, assume that our Course Allocation instance $J$ has a first-coalition-stable matching $M^{\prime}$. Then we take $M^{\prime}$, remove any pair containing a course of the form $c_{i}^{\prime}$, and swap courses and students for residents and hospitals to get a matching $M$ in the HRS instance $I$. This matching $M$ is also stable in $I$: if $(r_{j},h_{i})$ were to block $M$ in $I$, then either $(s_{i},c_{j})$ or $(s_{i},\{c_{j},c_{i}^{\prime}\})$ would block $M$ in $J$, a contradiction.
@@
Therefore, there exists a first-coalition-stable matching in $J$ if and only if there exists a stable matching in $I$. Thus, as HRS is NP-complete, so is ca-fc-find. ∎
@@ 4.2 Master lists
An important restriction of Course Allocation arises when students are ordered in a master list, for example by the average of their grades. Let ca-x-ml-find be the problem of finding a stable matching in a ca-ml instance $I$, for a given definition of stability x, or determining that none exists. When students are ranked according to a master list, we can use the Serial Dictatorship mechanism to find a stable matching, for any definition of stability. Serial Dictatorship is an algorithm that starts with an empty matching $M$ and, on round $i$, considers the $i$-th student in the master list, say $s_{i}$. Then, it considers every course on $s_{i}$’s preference list, from most to least preferred. For each such $c_{j}\in P(s_{i})$, it checks whether $c_{j}$ has some capacity left, whether adding $(s_{i},c_{j})$ to $M$ violates the credit limit of $s_{i}$, and, if there are downward-feasible constraints, whether adding $(s_{i},c_{j})$ to $M$ violates any of them. If it is possible to add $(s_{i},c_{j})$ to $M$, it does so, and proceeds to check the next course in $P(s_{i})$. This is clearly a $O(nm)$ time algorithm, where $n=|S|$ and $m=|C|$. We now show that Serial Dictatorship always finds a pair-stable (feasible) matching.
@@
In any ca-ml instance $I$, a pair-stable matching always exists, and one can use Serial Dictatorship to find such a matching in $O(nm)$ time, where $n=|S|$ and $m=|C|$.
@@
Aiming for a contradiction, assume that after executing Serial Dictatorship, student $s_{i}$ and course $c_{j}$ form a blocking pair. Then this means that either $c_{j}$ has some capacity left for $s_{i}$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ that is assigned to $c_{j}$. If the second case holds, $s_{i}\succ_{ML}s_{k}$, so $s_{i}$ got to choose courses before $s_{k}$, meaning that in the round of $s_{i}$, $s_{k}$ was not assigned to $c_{j}$ and $c_{j}$ had an open place that $s_{i}$ could have taken. If $s_{i}$ did not choose $c_{j}$ and instead chose other courses, then $s_{i}$ preferred her assigned courses over $c_{j}$, which contradicts the assumption that $s_{i}$ and $c_{j}$ form a blocking pair. The reasoning is similar if the first case holds instead. Therefore, the algorithm above always produces a pair-stable matching. ∎
By Proposition 1, a pair-stable matching is first-coalition-stable, coalition-stable, and pair-size-stable. Therefore, the matching found by Serial Dictatorship is first-coalition-stable, coalition-stable, and pair-size-stable, and thus ca-x-ml-find is solvable in polynomial time for any definition of stability x.
Moreover, the pair-stable matching found by Serial Dictatorship is the unique pair-stable matching. Assume otherwise, so that there are two different pair-stable matchings, $M$ and $M^{\prime}$, and let $s_{i}$ be the highest-ranked student such that $C_{M}(s_{i})\neq C_{M^{\prime}}(s_{i})$. That is, she is assigned different courses in each matching. Let $c_{j}$ be the most preferred course of $s_{i}$ in $C_{M}(s_{i})\ominus C_{M^{\prime}}(s_{i})$, the symmetric difference of $C_{M}(s_{i})$ and $C_{M^{\prime}}(s_{i})$. Without loss of generality, assume that $c_{j}\in C_{M}(s_{i})$. Then $(s_{i},c_{j})$ blocks $M^{\prime}$, because $s_{i}$ prefers to drop all courses in $C_{M^{\prime}}(s_{i})\setminus C_{M}(s_{i})$ and be allocated to $c_{j}$ instead, and the set $(C_{M}(s_{i})\cap C_{M^{\prime}}(s_{i}))\cup\{(s_{i},c_{j})\}$ is feasible, as it is a subset of the feasible set $C_{M}(s_{i})$. Meanwhile, because $s_{i}$ is the highest-ranked student that has different courses in $M$ and $M^{\prime}$, $c_{j}$ prefers $s_{i}$ over the students that it has in $M^{\prime}$. Hence, we obtain a contradiction to the pair stability of $M^{\prime}$, which implies that there is a unique pair-stable matching.

### v2

@@ 4.1 Arbitrary course preferences
Let us first consider the problem of finding a pair-size-stable matching. In order to do so, we use the following procedure. First, we group courses by number of credits, from largest to smallest. Let $\{O(c_{j}):c_{j}\in C\}=\{N_{1},\dots,N_{R}\}$ for some $R\geq 1$, where $N_{r}>N_{r+1}$ ($1\leq r\leq R-1$). Let $C_{r}=\{c_{j}\in C:O(c_{j})=N_{r}\}$ for each $r$ ($1\leq r\leq R$).
@@
0:  Course Allocation instance II
0:  return a pair-size-stable matching $M$ in II
@@
2:  for $r\in 1\dots R$ do
3:   while some student $s_{i}\in S$ satisfies $O(C_{M}(s_{i}))+N_{r}\leq T(s_{i})$ and $s_{i}$ has not applied to all feasible courses in $C_{r}$ on her list do
4:    $c_{j}:=$ $s_{i}$’s most-preferred feasible course in $C_{r}$ that $s_{i}$ has not yet applied to;
@@
It is clear that Algorithm 1 finds a matching $M$. Furthermore, if there are downward-feasible constraints, $M$ is feasible, because every $s_{i}$ applies only to courses $c_{j}$ such that adding $(s_{i},c_{j})$ to $M$ does not violate downward-feasible constraints. With appropriate data structures, the overall complexity of Algorithm 1 is $O(RL)=O(nm^{2})$, where $n=|S|$, $m=|C|$ and $L$ is the number of acceptable pairs. This is because for every student we consider each course in her preference list at most $m$ times, if there are downward-feasible constraints, as courses that were previously unfeasible might become feasible for some $s_{i}$ if she is rejected from some $c_{j}$. Meanwhile, when a student applies to a course, the course can perform all necessary steps in constant time, giving a total time complexity of $O(nm^{2})$, where $n=|S|$ and $m=|C|$. If there are no downward-feasible constraints then the time complexity is $O(nm)$. We now show that $M$ is pair-size-stable.
@@
[…]
Assume that the first case holds. The course $c_{j}$ is in subset $C_{r}$, meaning that it was considered in the $r$-th round. As $s_{i}$ has enough credits for $c_{j}$ after the execution of the algorithm, it had enough credits during round $r$, because matches between students and courses at the end of a given round $k$ cannot be broken subsequently, so after every round the number of remaining credits for $s_{i}$ stays the same or decreases. Thus, $s_{i}$ applied to $c_{j}$ during round $r$, so if they are not matched at the end of round $r$ then $c_{j}$ rejected $s_{i}$ for a better student. Hence, $(s_{i},c_{j})$ was not a size-blocking pair at the end of round $r$. Because all matches formed in round $r$ are part of the final pair-size-stable matching, $c_{j}$ prefers all of its students over $s_{i}$, so $(s_{i},c_{j})$ is not a size-blocking pair, a contradiction. Therefore, this case cannot occur.
Now, assume the second case holds, so $s_{i}$ prefers $c_{j}$ over every course $c_{k}$ from a subset of courses $D_{i}$ such that $s_{i}$ is assigned in $M$ to every $c_{k}\in D_{i}$ and $O(c_{j})\geq\sum_{c_{k}\in D_{i}}O(c_{k})$. As $O(c_{j})\geq\sum_{c_{k}\in D_{i}}O(c_{k})$, $O(c_{j})\geq O(c_{k})$ $\forall c_{k}\in D_{i}$, no $c_{k}\in D_{i}$ has a number of credits strictly greater than $O(c_{j})$. Additionally, no $c_{k}\in D_{i}$ has a number of credits equal to $O(c_{j})$, for if there were such a $c_{k}$ and $(s_{i},c_{j})$ was a size-blocking pair, then that would imply that in the round where $c_{j}$ and $c_{k}$ were considered, Deferred Acceptance did not find a stable matching, which is a contradiction. Therefore, $c_{j}$ has a number of credits strictly greater than every $c_{k}\in D_{i}$. However, this means that, during the algorithm’s execution, $c_{j}$ was considered in an earlier round than every $c_{k}\in D_{i}$, say round $r$. During that round, $s_{i}$ had enough remaining credits for $c_{j}$, because, after removing every course from $D_{i}$, it has enough credits for $c_{j}$ after the execution of the algorithm. Using a similar argument to the first case, $(s_{i},c_{j})$ was not a size-blocking pair at the end of round $r$, meaning that it is not a size-blocking pair upon termination of the algorithm, a contradiction. Therefore, this case also cannot occur. By assuming either case holds, we arrive at a contradiction, which is caused by the assumption that there exists a size-blocking pair. Therefore, no such pair exists, and the matching is pair-size-stable. ∎
@@
Figure 4: Instance of ca. This instance is derived from an hrs instance due to McDermid and Manlove McDermid and Manlove (2010).
Determining whether a pair-stable matching exists is NP-complete, which can be seen by considering a connection between ca and hrs (hrs was introduced in Section 1), which we now describe. With respect to the stability definition given by McDermid and Manlove McDermid and Manlove (2010), the problem of deciding whether a stable matching exists, given an instance of hrs, is NP-complete. Moreover this result holds even if each resident has size 1 or 2, each hospital has upper quota at most 2, and each preference list is of length at most 3. By replacing residents by courses and hospitals by students, we can then obtain a simple reduction to ca where each student $s_{i}$ satisfies $T(s_{i})\leq 2$ and each course $c_{j}$ satisfies $O(c_{j})\in\{1,2\}$ and $q(c_{j})=1$. Using this transformation, McDermid and Manlove’s stability definition in the hrs instance corresponds to pair stability in the transformed ca instance. The following result is therefore immediate.
Theorem 3 (McDermid and Manlove (2010)).
ca-p-find is NP-hard, even if there are no downward-feasible constraints, each student can take at most two credits, each course has one or two credits, each course has upper quota $1$, and each preference list is of length at most $3$.
Similarly, determining the existence of a first-coalition-stable matching is NP-hard, and we show that by transforming from hrs, though in this case a more complex reduction is required.
@@
ca-fc-find is NP-hard, even if there are no downward-feasible constraints, each student can take at most two credits, each course has one or two credits, each course has upper quota $1$, and each preference list is of length at most $4$.
@@
We prove this result by transforming from the problem of determining the existence of a stable matching in the hrs setting, which was shown to be NP-complete by McDermid and Manlove McDermid and Manlove (2010). In the hrs instance $I$ that we are transforming from, each resident has size $1$ or $2$, each hospital has upper quota $2$, and the length of every preference list is at most $3$.
@@
The transformation proceeds as follows. Every resident $r_{j}$ in $I$ becomes a course $c_{j}$ in the ca instance $J$, with upper quota $1$, and number of credits $1$ or $2$, depending on the size of $r_{j}$. Every hospital $h_{i}$ becomes a student $s_{i}$, with $T(s_{i})=2$. We use this bijection to transform every preference list. For example, if resident $r_{j}$ has preference list $h_{i}\succ_{r_{j}}h_{k}\succ_{r_{j}}h_{l}$, now course $c_{j}$ has preference list $s_{i}\succ_{c_{j}}s_{k}\succ_{c_{j}}s_{l}$. Additionally, for every student $s_{i}$ we create a dummy course $c_{i}^{\prime}$, with one credit, and append this course to the end of $s_{i}$’s preference list. The preference list of $c_{i}^{\prime}$ contains $s_{i}$ only.
@@
Assume that the hrs instance $I$ has a stable matching $M$. Then we build a first-coalition-stable matching $M^{\prime}$ in $J$ as follows. We take $M$ and swap residents and hospitals for courses and students, respectively. Then, for every student $s_{i}$ such that $O(C_{M^{\prime}}(s_{i}))<2$, we add the pair $(s_{i},c_{i}^{\prime})$ to $M^{\prime}$. We claim that $M^{\prime}$ is first-coalition-stable. Assume $M^{\prime}$ is not, which implies the existence of a first-blocking coalition of $M^{\prime}$ in $J$ involving one student $s_{i}$. Let us go through each possibility:
@@
  • Student $s_{i}$ has only one credit in $M^{\prime}$ (which must be from course $c_{i}^{\prime}$, according to the construction of $M^{\prime}$), and forms a first-blocking coalition with a course $c_{j}$, with one or two credits. In this case, hospital $h_{i}$ would have no residents in $M$, and resident $r_{j}$ is unmatched in $M$ or prefers $h_{i}$ to $M(r_{j})$, so $(r_{j},h_{i})$ blocks $M$ in $I$, a contradiction.
@@
  • Student $s_{i}$ has two credits in $M^{\prime}$, one of which comes from $c_{i}^{\prime}$ and the other from some $c_{k}$, and she forms a first-blocking coalition with course $c_{j}$ (and possibly with one other course also). As $c_{i}^{\prime}$ is the least-preferred course on $s_{i}$’s preference list, it follows that $s_{i}$ prefers $c_{j}$ over $c_{i}^{\prime}$. Suppose firstly that $c_{j}$ has one credit. As $C_{M^{\prime}}(s_{i})=\{c_{k},c_{i}^{\prime}\}$, $(r_{k},h_{i})\in M$ where $r_{k}$ has size $1$. Moreover $h_{i}$ has room for resident $r_{j}$, also of size $1$. Resident $r_{j}$ is unmatched in $M$ or prefers $h_{i}$ over $M(r_{j})$, and hence $(r_{j},h_{i})$ forms a blocking pair of $M$ in $I$, a contradiction. If $c_{j}$ has two credits, then $s_{i}$ prefers $c_{j}$ over $c_{k}$, which implies that in $I$ hospital $h_{i}$ prefers resident $r_{j}$ over its assigned resident $r_{k}$, and $r_{j}$ is unmatched in $M$ or prefers $h_{i}$ to $M(r_{j})$, so $(r_{j},h_{i})$ blocks $M$ in $I$, a contradiction.
@@
  • Student $s_{i}$ has two credits in $M^{\prime}$, which come from two courses with one credit, namely $c_{p}$ and $c_{q}$ (neither of which is equal to $c_{i}^{\prime}$), where $c_{p}\succ_{s_{i}}c_{q}$, or one course with two credits, namely $c_{p}$. Suppose that $s_{i}$ forms a first-blocking coalition with either one course $c_{j}$, or with two courses, $c_{j}$ and $c_{k}$, such that $c_{j}\succ_{s_{i}}c_{k}$. Then, in $I$, hospital $h_{i}$ prefers resident $r_{j}$ over either resident $r_{p}$ or residents $r_{p}$ and $r_{q}$, and $r_{j}$ is unmatched in $M$ or prefers $h_{i}$ to $M(r_{j})$, so $(r_{j},h_{i})$ blocks $M$ in $I$, a contradiction.
@@
Similarly, assume that the ca instance $J$ has a first-coalition-stable matching $M^{\prime}$. Then we take $M^{\prime}$, remove any pair containing a course of the form $c_{i}^{\prime}$, and swap courses and students for residents and hospitals to obtain a matching $M$ in the hrs instance $I$. This matching $M$ is also stable in $I$: if $(r_{j},h_{i})$ were to block $M$ in $I$, then either $(s_{i},c_{j})$ or $(s_{i},\{c_{j},c_{i}^{\prime}\})$ would form a first-blocking coalition of $M$ in $J$, a contradiction.
@@
Therefore, there exists a first-coalition-stable matching in $J$ if and only if there exists a stable matching in $I$. Thus, as hrs is NP-complete, ca-fc-find is NP-hard. ∎
@@ 4.2 Master lists
An important restriction of ca arises when students are ordered in a master list, for example by Grade Point Average (GPA). Let ca-x-ml-find be the problem of finding a stable matching in a ca-ml instance $I$, for a given definition of stability x, or determining that none exists. When students are ranked according to a master list, we can use a modification of the Serial Dictatorship mechanism to find a stable matching, for any definition of stability. The algorithm starts with an empty matching $M$ and, on round $i$, considers the $i$-th student in the master list, say $s_{i}$. Then, it considers every course on $s_{i}$’s preference list, from most to least preferred. For each such $c_{j}\in P(s_{i})$, it checks whether $c_{j}$ has some capacity left, whether adding $(s_{i},c_{j})$ to $M$ does not violate the credit limit of $s_{i}$, and, if there are downward-feasible constraints, whether adding $(s_{i},c_{j})$ to $M$ does not violate any of them. If it is possible to add $(s_{i},c_{j})$ to $M$, it does so, and proceeds to check the next course in $P(s_{i})$. This algorithm is described in pseudocode form in Algorithm 2. This is clearly an $O(L)$ time algorithm, where $L$ is the total number of acceptable pairs. We now show that Algorithm 2 always finds a pair-stable (feasible) matching.
Algorithm 2 Algorithm to find an x-stable (feasible) matching in an instance of ca-x-ml-find
1:  $M:=\emptyset$;
2:  for $s_{i}$ in $\succ_{MLS}$ in decreasing order of preference do
3:   $p:=1$;      {$p$ points to first course in $s_{i}$’s list}
4:   while $O(C_{M}(s_{i}))<T(s_{i})$ and $p\leq|P(s_{i})|$ do
5:    $c_{j}:=$ course at position $p$ of $\succ_{s_{i}}$;
6:    if $O(C_{M}(s_{i}))+O(c_{j})\leq T(s_{i})$ and $|S_{M}(c_{j})|<q^{+}(c_{j})$ and $C_{M}(s_{i})\cup\{c_{j}\}$ is feasible then
7:     $M:=M\cup\{(s_{i},c_{j})\}$;
8:    end if
9:    pp++;
10:   end while;
11:  end for
12:  return $M$;
@@
In any ca-ml instance $I$, a pair-stable (feasible) matching always exists, and one can use Algorithm 2 to find such a matching in $O(L)$ time, where $L$ is the total number of acceptable pairs.
@@
Aiming for a contradiction, assume that after executing Algorithm 2, student $s_{i}$ and course $c_{j}$ form a blocking pair. Then this means that either $c_{j}$ has some capacity left for $s_{i}$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ that is assigned to $c_{j}$. If the second case holds, $s_{i}\succ_{MLS}s_{k}$, so $s_{i}$ got to choose courses before $s_{k}$, meaning that in the round of $s_{i}$, $s_{k}$ was not assigned to $c_{j}$ and $c_{j}$ had an open place that $s_{i}$ could have taken. If $s_{i}$ did not choose $c_{j}$ and instead chose other courses, then $s_{i}$ would prefer her assigned courses over $c_{j}$, which contradicts the assumption that $s_{i}$ and $c_{j}$ form a blocking pair. The reasoning is similar if the first case holds instead. Therefore, the algorithm above always produces a pair-stable matching. ∎
By Proposition 1, a pair-stable matching is first-coalition-stable, coalition-stable, and pair-size-stable. Therefore, the matching found by Algorithm 2 is first-coalition-stable, coalition-stable, and pair-size-stable, and thus ca-x-ml-find is solvable in polynomial time for any definition of stability x.
Moreover, the pair-stable (feasible) matching found by Algorithm 2 is the unique pair-stable (feasible) matching. Assume otherwise, so that there are two different pair-stable (feasible) matchings, namely $M$ and $M^{\prime}$, and let $s_{i}$ be the highest-ranked student such that $C_{M}(s_{i})\neq C_{M^{\prime}}(s_{i})$. That is, she is assigned a different set of courses in each matching. Let $c_{j}$ be the most-preferred course of $s_{i}$ in $C_{M}(s_{i})\oplus C_{M^{\prime}}(s_{i})$, the symmetric difference of $C_{M}(s_{i})$ and $C_{M^{\prime}}(s_{i})$. Without loss of generality, assume that $c_{j}\in C_{M}(s_{i})$. Then $(s_{i},c_{j})$ blocks $M^{\prime}$, because $s_{i}$ prefers to drop all courses in $C_{M^{\prime}}(s_{i})\setminus C_{M}(s_{i})$ and be allocated to $c_{j}$ instead; this is the case because $O(C_{M}(s_{i}))\leq T(s_{i})$ and $c_{j}\in C_{M}(s_{i})\setminus C_{M^{\prime}}(s_{i})$ implies that $O(C_{M}(s_{i})\cap C_{M^{\prime}}(s_{i}))+O(c_{j})\leq T(s_{i})$; also the set $(C_{M}(s_{i})\cap C_{M^{\prime}}(s_{i}))\cup\{(s_{i},c_{j})\}$ is feasible, as it is a subset of the feasible set $C_{M}(s_{i})$; finally, because $s_{i}$ is the highest-ranked student that has different courses in $M$ and $M^{\prime}$, $c_{j}$ prefers $s_{i}$ over the students that it has in $M^{\prime}$ but not in $M$. Hence, we obtain a contradiction to the pair stability of $M^{\prime}$, which implies that there is a unique pair-stable matching.